Project

General

Profile

NDOS Commissioning Meeting 2011/05/16

Attendance: Jon, Mat, Peter, Mark Mathis, Mark Messier, Susan, Rick K., Greg, Denis, Minerba, Sue, Jarek, Leon, Gavin

Agenda:

  1. NDOS Status and Channel Recovery Plans
  2. DAQ/Electronics Update & Plans
  3. DCS Update & Plans
  4. AOB (All)

NDOS Status and Channel Recovery Plans

  • One phone call to DAQ group over the weekend on Sunday evening, regarding DAQAppMgr. Kurt is investigating.
  • Harvard responded to Mat's request for ribbon cables, they should be arriving very soon.
  • Peter reported that the NuMI target He leak has worsened, so the switch to HE beam run may happen earlier than expected.
  • Mat: dehumidifier in NDSB was found to be on, but set to 60%. Mat lowered it to 30%, and it seems to have improved the situation.

DAQ/Electronics Update & Plans

  • DAQ stress tests have begun today.
  • So far they have found that a trigger rate of ~100Hz causes sporadic triggers.
  • DDS upgrade over the weekend should improve stability

DCS Update & Plans

  • Peter reported that tests on detector showed some instability
  • Sue put a note out on the baseline tests performed on the teststand; timing glitches were seen even when FEBs were not being read out.

AOB (All)